OpenStudy (anonymous):

solve the system of equations by using the substitution method 3x-5y=8 y=x-2

OpenStudy (saifoo.khan):

Who's gonna win?

OpenStudy (anonymous):

how does coolsector do that, does he have stealth technology?

OpenStudy (anonymous):

omg i deleted it without coping :( y =x -2 plug it into the second : 3x - 5(x-2) = 8 -2x +10 = 8 -2x = -2 x= 1 plug it into y = x- 2 y =1-2 = -1 (1,-1)

OpenStudy (anonymous):

i wanted to give agdg.... a medal so i was waiting :})

hero (hero):

y = 8/5 y = x-2 8/5 - 10/5 = x -2/5 = x

OpenStudy (anonymous):

see! she's always givning me medal i wanted to reciprocate!

hero (hero):

Ooops 8/5 + 10/5 = 18/5 x = 18/5

OpenStudy (anonymous):

hero, seriously i never see coolsector then his posts appear how does he do it

hero (hero):

I saw him

OpenStudy (anonymous):

wierd server stuff i guess